OpenStudy (anonymous):

How to complete the square such that ax^2 + bx + c = a(x-p)^2 + q?

OpenStudy (anonymous):

Do I do it like this: Please check my answers.. Work out the multiplication on the (x-p)^2 part. y = a(x-p)^2 + q y = a(x^2 - 2px + p^2) + q y = ax^2 - 2apx + ap^2 + q So now we have two formulas, y = ax^2 + bx + c y = ax^2 - 2apx + (ap^2 + q) They're equal to each other for all values of x, so the coefficients must be the same. The ax^2 is the same in both. a=a bx = -2apx so b = -2ap and p = -b/2a c = ap^2 + q so q = c - ap^2 since p = -b/2a, you can also say q = c - a(-b/2a)^2 q = c - a(b^2/4a^2) q = c - b^2/4a

OpenStudy (lalaly):

\[=a(x^2+\frac{b}{a}x+\frac{c}{a})\]\[=a[(x+\frac{b}{2a})^2-(\frac{b}{2a})^2+\frac{c}{a}]\]\[ p=\frac{-b}{2a}\]\[q=-(\frac{b}{2a})^2+\frac{c}{a}\]